OpenStudy (anonymous):

–(–33 + 18) = 33 + (–18)

OpenStudy (anonymous):

What property is shown by the equation?

OpenStudy (anonymous):

@KayleeMullins

OpenStudy (anonymous):

distributive (property)

OpenStudy (anonymous):

commutative property

OpenStudy (anonymous):

A. commutative property B. associative property C. opposite of a sum property

OpenStudy (anonymous):

these are the options

OpenStudy (phi):

In English , "distribute" means hand out to each person in math, think of the -1 multiplying "out front" being "handed out" to each term inside the parens

OpenStudy (anonymous):

?

OpenStudy (anonymous):

can you help @KayleeMullins

OpenStudy (anonymous):

its opposite of a sum property :)

OpenStudy (anonymous):

thx I have more

OpenStudy (anonymous):

yeah :)

OpenStudy (anonymous):

–10 + (–15) = –15 + (–10)

OpenStudy (anonymous):

Comunative property :) @shadow2239

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

4 + (29 + 16) = 4 + (16 + 29)

OpenStudy (anonymous):

Commutative property

OpenStudy (anonymous):

thx one more

OpenStudy (anonymous):

–12 + (12 + 27) = (–12 + 12) + 27

OpenStudy (anonymous):

Associative Property :)

OpenStudy (anonymous):

thx so much